Find two consecutive positive intergers such that the sum of their square is 85.

Answer by checkley71(8403)

X^2+(X+1)^2=85
X^2+X^2+2X+1=85
2X^2+2X+1-85=0
2X^2+2X-84=0
2(X^2+X-42)
2(X+7)(X-6)=0
X+7=0
X=-7 ANSWER.
X-6=0
X=6 ANSWER.
PROOF
6^2+7^2=85
36+49=85
85=85